express in terms of theta: csc (90 degrees + theta) sin (180 degrees + theta) - tan (270 degrees - theta).
Answer by venugopalramana(3286)

express in terms of theta: csc (90 degrees + theta) sin (180 degrees + theta) - tan (270 degrees - theta).
LET US USE A INSTEAD OF THETA.....USING ALL SILVER TEA CUPS FOR SIGN
=-SEC(A)*(-SIN(A))*(COT(A))=(1/COS(A))(SIN(A))(COS(A)/SIN(A))=1